# Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation

### v2

@@ 4.3 Model Robustness
Table 3: Adversarial Robustness of Merged Models on the SST2 and CoLA Datasets when Executing StressTest prompt attack method. The best and second-best results are marked in bold and underlined fonts.
We employ three Prompt Attack Methods supported by the promptbench codebase (DeepWordBug, BERTAttack, and StressTest) (Zhu et al., 2024) to evaluate the adversarial robustness of three merged models (Math & Code, LM & Math, and LM & Code) obtained through the task arithmetic method. To balance experimental effectiveness with computational efficiency, we randomly selected the positions of the three attacked words in the prompts when executing the DeepWordBug and BERTAttack attacks. Adversarial robustness is assessed using the Performance Drop Rate (PDR) (Zhu et al., 2023), where a lower PDR indicates stronger robustness. Further details on PDR can be found in Appendix D.
As shown in Table 3, M3 improves the adversarial robustness of the merged models in most cases with the StressTest Prompt Attack Method. For example, with M3, the PDR of Math & Code decreased by 3.2% on the SST2 dataset and by 92.12% on the CoLA dataset, while the PDR of LM & Code decreased by 30.36% on SST2 and by 15.75% on CoLA. Furthermore, in most cases, M3 not only improves the adversarial robustness of the merged models but also enhances their performance metrics (accuracy and MCC) on the SST2 and CoLA datasets. Specifically, with M3, Math & Code demonstrates a 28.56% increase in accuracy on SST2 and a 26.18% increase in MCC on CoLA, while LM & Code achieves a 62.62% increase in accuracy on SST2. These results show that M3 effectively enhances both the adversarial robustness and the performance of the merged models in sentiment analysis and grammar correctness tasks. Detailed experimental results for the remaining Prompt Attack Methods (DeepWordBug and BERTAttack) are presented in Appendix B.
@@ 4.4 Comparison of Mixup Model Merge and DARE
DARE is a model sparsification method proposed by (Yu et al., 2024), with a more detailed introduction provided in Appendix E. This section takes DARE as an example to discuss the relationship between the existing model sparsification method and M3. We integrated M3 and DARE with three existing model merging techniques—Average Merging, Task Arithmetic, and TIES-Merging—to compare these two plug-and-play methods. Detailed experimental results are provided in Appendix F. Additionally, in the DARE method, the drop rate hyperparameter is set to 0.2.
From Table 8, it can be concluded that M3 generally outperforms DARE, exhibiting significantly superior performance on certain datasets. For instance, the Math & Code model achieves a pass@1 score of 9.8% on the MBPP dataset when combined with DARE, while this score increases to 19% when combined with M3.
As shown in Table 8, Combining DARE and M3 generally results in better model merging performance. This suggests that M3 and DARE can complement each other. Moreover, in some cases, M3 alone can deliver the best results, while DARE alone only achieves the best performance in very few cases, further demonstrating the superiority of M3.
@@ 5 Conclusion
Addressing the issue that existing methods overlook the varying contribution ratios of different models, we propose Mixup Model Merge (M3), a novel approach based on the randomized linear interpolation strategy of Mixup. M3 performs linear interpolation in the parameter space between two task-specific large language models, with interpolation coefficients sampled from a Beta distribution to explore a diverse range of contribution ratios. This method leverages controllable randomness to significantly outperform traditional equal-ratio merging. Extensive experiments demonstrate that M3 not only improves merged model performance across multiple tasks but also enhances out-of-distribution and adversarial robustness, and effectively combines with model sparsification methods such as DARE. By tuning the Beta distribution parameters, M3 balances efficiency and diversity in exploring contribution ratios.
6 Limitations
Despite its advantages, the M3 method has several limitations: (1) Our method is currently limited to pairwise model merging, as random linear interpolation is typically defined between two entities. Extending it to merge three or more models requires new strategies for sampling multi-way interpolation coefficients, which introduces additional complexity and remains an open direction for future work. (2) This study focuses on merging homogeneous models fine-tuned from the same pretrained backbone. Applying M3 to heterogeneous models with different architectures or fine-tuning paths poses new challenges and will be explored in future work. (3) Our method may also be extended to broader applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) models to reduce the alignment tax (FINE-TUNING, ).
Appendix A Detailed Experimental Settings
A.1 Task-Specific Fine-Tuned LLMs and Datasets Details
@@
  • WizardLM-13B is an instruction-following model based on Llama-2-13b, designed to improve open-domain instruction-following. Using the Evol-Instruct method (Xu et al., 2024), it generates high-complexity instruction data to reduce human annotation and enhance generalization. The model undergoes supervised fine-tuning with AI-generated data, followed by refinement via RLHF. Evaluation results show that Evol-Instruct-generated instructions outperform human-written ones, and WizardLM-13B surpasses ChatGPT in high-complexity tasks. In GPT-4 automated evaluation, it achieves over 90% of ChatGPT’s performance in 17 out of 29 tasks, demonstrating the effectiveness of AI-evolved instruction fine-tuning (Xu et al., 2024).
  • WizardMath-13B, optimized from Llama-2-13b, is designed for mathematical reasoning and enhances Chain-of-Thought (CoT) (Wei et al., 2022) capabilities. It uses Reinforcement Learning from Evol-Instruct Feedback to evolve math tasks and improve reasoning. Trained on GSM8K and MATH datasets, it excels in both basic and advanced math problems. In evaluations, WizardMath-Mistral 7B outperforms all open-source models with fewer training data, while WizardMath 70B surpasses GPT-3.5-Turbo, Claude 2, and even early GPT-4 versions in mathematical reasoning tasks.
  • llama-2-13b-code-alpaca is a code generation model fine-tuned from Llama-2-13b, designed to enhance code understanding and generation. It follows the same training approach as Stanford Alpaca (Taori et al., 2023) but focuses on code-related tasks. The model is fine-tuned with 20K instruction-following code samples generated using the Self-Instruct method (Wang et al., 2022). However, as it has not undergone safety fine-tuning, caution is required when using it in production environments.
@@
  • AlpacaEval (Li et al., 2023) is an LLM-based automated evaluation metric that assesses model performance by testing on a fixed set of 805 instructions and computing the win rate of the evaluated model against a baseline. The evaluation process involves an LLM-based evaluator that compares the responses and determines the probability of preferring the evaluated model. In this paper, we use AlpacaEval 2.0 (Dubois et al., 2024). To reduce costs, we use chatgpt_fn for evaluation.
@@
  • GSM8K is a dataset of 8.5K high-quality, linguistically diverse grade school math word problems, designed to evaluate the multi-step mathematical reasoning abilities of large language models. It consists of 7.5K training problems and 1K test problems. In this paper, we use the 1K test set for evaluation (Cobbe et al., 2021).
  • MATH is a dataset containing 12,500 competition-level math problems, designed to evaluate and enhance the problem-solving abilities of machine learning models. It consists of 7,500 training problems and 5,000 test problems. We use the 5,000 test set for evaluation (Hendrycks et al., 2021).
@@
  • HumanEval is a dataset consisting of 164 hand-written programming problems, designed to evaluate the functional correctness of code generation models. Each problem includes a function signature, docstring, function body, and unit tests. The dataset tests models’ language comprehension, reasoning, and algorithmic abilities (Chen et al., 2021).
  • MBPP is a dataset containing 974 programming problems designed to evaluate a model’s ability to synthesize Python programs from natural language descriptions. The problems range from basic numerical operations to more complex tasks involving list and string processing. The test set consists of 500 problems, which are used for evaluation in this paper (Austin et al., 2021).
A.2 Hyperparameter Setting Details in Model Merging Methods
@@
Merging Method Model Hyperparameter Values
TIES-Merging LM & Math scaling_term=0.5, retain_ratio=0.5
LM & Code scaling_term=1.0, retain_ratio=0.7
Math & Code scaling_term=1.0, retain_ratio=0.5
@@
A.3 Out-of-Distribution Dataset Selection Details
LiveBench (White et al., 2024) is a dynamic benchmark for large language models, featuring frequently updated questions and diverse tasks. To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench, respectively, deliberately avoiding the fine-tuning domains of the merged fine-tuned models. These tasks were released after November 2023, whereas WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca were all introduced earlier. Furthermore, their shared Llama-2-13b backbone was trained on data only up to July 2023. Consequently, these factors collectively ensure that the evaluation remains OOD in the temporal dimension.
Finally, we acknowledge the limitations of these datasets. For large models like Llama-2-13b, identifying truly OOD datasets is difficult, as their training data likely covers similar distributions. These datasets are better described as "out-of-example", representing instances not explicitly seen during training. As discussed in (Wang et al., 2023), distribution shifts can occur across domains and time. While Llama-2-13b may have been trained on datasets for tasks like instruction-following, coding, and language comprehension, the datasets we selected remain valuable for OOD evaluation by capturing temporal shifts, providing insights into robustness over time.
A.4 Adversarial Robustness Evaluation Experiments Setting Details
PromptBench (Zhu et al., 2024) is a unified library designed for evaluating LLMs, providing a standardized and extensible framework. It includes several key components such as prompt construction, prompt engineering, dataset and model loading, adversarial prompt attacks, dynamic evaluation protocols, and analysis tools.
We use the Adversarial Prompt Attacks module in PromptBench aims to evaluate the robustness of LLMs against adversarial prompts. We employ three methods to perform adversarial attacks on prompts to evaluate the adversarial robustness of the merged models: DeepWordBug (Gao et al., 2018), BERTAttack (Li et al., 2020), and StressTest (Naik et al., 2018), representing Character-level, Word-level, and Sentence-level attacks, respectively.
(a) The operational steps of TIES-Merging.
The evaluation is conducted on the Sentiment Analysis dataset (SST2 (Socher et al., 2013)) and the Grammar Correctness dataset (CoLA (Warstadt, 2019)):
  • SST2 (Socher et al., 2013): A sentiment analysis dataset designed to assess whether a given sentence conveys a positive or negative sentiment.
  • CoLA (Warstadt, 2019): A dataset for grammar correctness, where the model must determine whether a sentence is grammatically acceptable.
Appendix B Additional Experimental Results on Adversarial Robustness
@@
Appendix C Integrating M3 into the TIES-Merging Model Merging Method

### v3

@@ 4.1 Experimental Setup
Table 2: Adversarial robustness of merged models on SST2 and CoLA with StressTest prompt attacks. Best and second-best results are marked in bold and underlined, respectively.
Merging
Methods
Models
Use M3
Use
DARE
Instruction
Following
Mathematical
Reasoning
Code Generating
AlpacaEval GSM8K MATH HumanEval MBPP
Average
Merging
LM
& Math
Yes No 44.40 66.26 13.80 - -
No Yes 44.22 66.57 12.96 - -
Yes Yes 43.53 66.57 14.12 - -
LM
& Code
Yes No 43.91 - - 37.20 34.40
No Yes 38.81 - - 31.71 32.40
Yes Yes 40.31 - - 36.59 37.00
Math
& Code
[…]
No Yes - 56.18 10.28 6.10 7.80
Yes Yes - 64.97 13.54 9.76 21.20
Task
Arithmetic
LM
& Math
Yes No 41.65 66.34 13.74 - -
No Yes 49.00 66.64 13.02 - -
Yes Yes 44.90 67.32 13.74 - -
LM
& Code
Yes No 46.64 - - 35.37 33.80
No Yes 41.47 - - 35.98 33.00
Yes Yes 45.20 - - 35.98 35.20
Math
& Code
Yes No - 63.53 13.94 7.93 19.00
No Yes - 65.05 13.96 10.37 9.80
Yes Yes - 65.13 14.32 8.54 18.00
TIES-
Merging
LM
& Math
Yes No 38.73 18.57 2.48 - -
No Yes 37.92 18.04 2.34 - -
Yes Yes 39.93 19.26 2.82 - -
LM
& Code
Yes No 44.96 - - 25.61 30.80
No Yes 43.13 - - 0.0 0.0
Yes Yes 45.65 - - 26.83 33.20
Math
& Code
Yes No - 64.75 14.16 9.76 21.4
No Yes - 64.82 13.88 10.37 23.60
Yes Yes - 64.75 14.78 9.15 19.60
4.2 Merging Task-Specific Fine-Tuned LLMs
Performance Gains of M3 Across Merging Methods and Tasks.
In addition, we also compare our proposed method M3 with existing model merging methods. The details are in Appendix E.
Mitigating Suboptimal Fine-Tuning with M3.
4.3 Model Robustness
Out-of-Distribution Robustness.
To ensure the evaluation datasets reflect true OOD scenarios, we select recently released datasets from domains not seen during fine-tuning. Accordingly, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing. As shown in Figure 4, M3 consistently improves OOD performance across all model pairs and merging methods. For example, under Task Arithmetic, M3 yields gains of 1.9%, 1.6%, and 6% on the three respective tasks. Similar improvements are observed with Average Merging (1.5%, 0.7%, 4%) and TIES-Merging (1.1%, 0.6%, 14%). These results highlight M3’s effectiveness in enhancing OOD generalization.
Adversarial robustness.
We evaluate the adversarial robustness of three merged models (Math & Code, LM & Math, and LM & Code) using three prompt attack methods from PromptBench (Zhu et al. 2024): DeepWordBug, BERTAttack, and StressTest. For DeepWordBug and BERTAttack, we randomly select three word positions per prompt. Robustness is measured by Performance Drop Rate (PDR) (Zhu et al. 2023), with lower PDR indicating stronger robustness (see Appendix F for details). As shown in Table 2, M3 notably improves robustness under StressTest attacks. For instance, it reduces PDR by 3.2% (SST2) and 92.12% (CoLA) for Math & Code, and by 30.36% (SST2) and 15.75% (CoLA) for LM & Code. Additionally, M3 improves performance metrics such as accuracy and MCC—e.g., accuracy for LM & Code on SST2 increases by 62.62%. These results indicate that M3 enhances both robustness and task performance. Results for DeepWordBug and BERTAttack are provided in Appendix G.
@@ 4.4 Comparison of Mixup Model Merge and DARE
DARE is a model sparsification method proposed by Yu et al. (2024) (see Appendix H for details). We compare M3 and DARE by integrating both into three representative model merging methods: Average Merging, Task Arithmetic, and TIES-Merging. The drop rate for DARE is fixed at 0.2. As shown in Table 3, M3 generally outperforms DARE and achieves significantly better results on several datasets. For example, on MBPP, the pass@1 score of the Math & Code model increases from 9.8% with DARE to 19% with M3. Moreover, combining M3 with DARE typically leads to even stronger performance, suggesting they are complementary. Notably, M3 alone often yields the best results, while DARE alone rarely does, further highlighting the effectiveness of M3.
@@ 5 Conclusion
Addressing the issue that existing methods overlook the varying contribution ratios of different models, we propose Mixup Model Merge (M3), a novel approach based on the randomized linear interpolation strategy of Mixup. M3 performs linear interpolation in the parameter space between two task-specific large language models, with interpolation coefficients sampled from a Beta distribution to explore a diverse range of contribution ratios. This method leverages controllable randomness to significantly outperform traditional equal-ratio merging. Extensive experiments demonstrate that M3 not only improves merged model performance across multiple tasks but also enhances out-of-distribution and adversarial robustness, and effectively combines with model sparsification methods such as DARE. By tuning the Beta distribution parameters, M3 balances efficiency and diversity in exploring contribution ratios. This simple and efficient method offers new insights for optimizing model merging strategies.
In the future, M3 may also be extended to broader applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) models (Ouyang et al. 2022), potentially reducing the alignment tax associated with post-hoc alignment techniques like fine-tuning.
Appendix A Task-Specific Fine-Tuned LLMs and Datasets Details
@@
  • WizardLM-13B is an instruction-following model based on Llama-2-13b, designed to improve open-domain instruction-following. Using the Evol-Instruct method (Xu et al. 2024), it generates high-complexity instruction data to reduce human annotation and enhance generalization. The model undergoes supervised fine-tuning with AI-generated data, followed by refinement via RLHF. Evaluation results show that Evol-Instruct-generated instructions outperform human-written ones, and WizardLM-13B surpasses ChatGPT in high-complexity tasks. In GPT-4 automated evaluation, it achieves over 90% of ChatGPT’s performance in 17 out of 29 tasks, demonstrating the effectiveness of AI-evolved instruction fine-tuning (Xu et al. 2024).
  • WizardMath-13B, optimized from Llama-2-13b, is designed for mathematical reasoning and enhances Chain-of-Thought (CoT) (Wei et al. 2022) capabilities. It uses Reinforcement Learning from Evol-Instruct Feedback to evolve math tasks and improve reasoning. Trained on GSM8K and MATH datasets, it excels in both basic and advanced math problems. In evaluations, WizardMath-Mistral 7B outperforms all open-source models with fewer training data, while WizardMath 70B surpasses GPT-3.5-Turbo, Claude 2, and even early GPT-4 versions in mathematical reasoning tasks.
  • llama-2-13b-code-alpaca is a code generation model fine-tuned from Llama-2-13b, designed to enhance code understanding and generation. It follows the same training approach as Stanford Alpaca (Taori et al. 2023) but focuses on code-related tasks. The model is fine-tuned with 20K instruction-following code samples generated using the Self-Instruct method (Wang et al. 2022). However, as it has not undergone safety fine-tuning, caution is required when using it in production environments.
@@
  • AlpacaEval (Li et al. 2023) is an LLM-based automated evaluation metric that assesses model performance by testing on a fixed set of 805 instructions and computing the win rate of the evaluated model against a baseline. The evaluation process involves an LLM-based evaluator that compares the responses and determines the probability of preferring the evaluated model. In this paper, we use AlpacaEval 2.0 (Dubois et al. 2024). To reduce costs, we use chatgpt_fn for evaluation.
@@
  • GSM8K is a dataset of 8.5K high-quality, linguistically diverse grade school math word problems, designed to evaluate the multi-step mathematical reasoning abilities of large language models. It consists of 7.5K training problems and 1K test problems. In this paper, we use the 1K test set for evaluation (Cobbe et al. 2021).
  • MATH is a dataset containing 12,500 competition-level math problems, designed to evaluate and enhance the problem-solving abilities of machine learning models. It consists of 7,500 training problems and 5,000 test problems. We use the 5,000 test set for evaluation (Hendrycks et al. 2021).
@@
  • HumanEval is a dataset consisting of 164 hand-written programming problems, designed to evaluate the functional correctness of code generation models. Each problem includes a function signature, docstring, function body, and unit tests. The dataset tests models’ language comprehension, reasoning, and algorithmic abilities (Chen et al. 2021).
  • MBPP is a dataset containing 974 programming problems designed to evaluate a model’s ability to synthesize Python programs from natural language descriptions. The problems range from basic numerical operations to more complex tasks involving list and string processing. The test set consists of 500 problems, which are used for evaluation in this paper (Austin et al. 2021).
Appendix B Out-of-Distribution Dataset Selection Details
LiveBench (White et al. 2024) is a dynamic benchmark for large language models, featuring frequently updated questions and diverse tasks. To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench, respectively, deliberately avoiding the fine-tuning domains of the merged fine-tuned models. These tasks were released after November 2023, whereas WizardLM-13B, WizardMath-13B, and llama-2-13b-code-alpaca were all introduced earlier. Furthermore, their shared Llama-2-13b backbone was trained on data only up to July 2023. Consequently, these factors collectively ensure that the evaluation remains OOD in the temporal dimension.
Finally, we acknowledge the limitations of these datasets. For large models like Llama-2-13b, identifying truly OOD datasets is difficult, as their training data likely covers similar distributions. These datasets are better described as ”out-of-example”, representing instances not explicitly seen during training. As discussed in (Wang et al. 2023), distribution shifts can occur across domains and time. While Llama-2-13b may have been trained on datasets for tasks like instruction-following, coding, and language comprehension, the datasets we selected remain valuable for OOD evaluation by capturing temporal shifts, providing insights into robustness over time.
Appendix C Adversarial Robustness Evaluation Experiments Setting Details
PromptBench (Zhu et al. 2024) is a unified library designed for evaluating LLMs, providing a standardized and extensible framework. It includes several key components such as prompt construction, prompt engineering, dataset and model loading, adversarial prompt attacks, dynamic evaluation protocols, and analysis tools.
We use the Adversarial Prompt Attacks module in PromptBench aims to evaluate the robustness of LLMs against adversarial prompts. We employ three methods to perform adversarial attacks on prompts to evaluate the adversarial robustness of the merged models: DeepWordBug (Gao et al. 2018), BERTAttack (Li et al. 2020), and StressTest (Naik et al. 2018), representing Character-level, Word-level, and Sentence-level attacks, respectively.
(a) The operational steps of TIES-Merging.
The evaluation is conducted on the Sentiment Analysis dataset (SST2 (Socher et al. 2013)) and the Grammar Correctness dataset (CoLA (Warstadt 2019)):
  • SST2 (Socher et al. 2013): A sentiment analysis dataset designed to assess whether a given sentence conveys a positive or negative sentiment.
  • CoLA (Warstadt 2019): A dataset for grammar correctness, where the model must determine whether a sentence is grammatically acceptable.
Appendix D Hyperparameter Setting Details in Model Merging Methods
@@
Merging Method
Model
Hyperparameter Values
TIES-Merging
LM & Math
scaling_term=0.5, retain_ratio=0.5
LM & Code
scaling_term=1.0, retain_ratio=0.7
Math & Code
scaling_term=1.0, retain_ratio=0.5
@@
Appendix E Comparison of M3 and Existing Model Merging Approaches
Models
Merging Methods
Instruction
Following
Mathematical
Reasoning
Code
Generating
Avg.
AlpacaEval GSM8K MBPP
LM
& Math
Average Merging
45.28 66.34 - 55.81
Task Arithmetic
45.78 66.34 - 56.06
TIES-Merging
38.63 14.56 - 26.60
DARE 49.00 66.64 - 57.82
DELLA 44.03 48.45 - 46.24
M3 (Ours) 44.90 67.32 - 56.11
LM
& Code
Average Merging
36.60 - 32.00 34.3
Task Arithmetic
44.64 - 33.60 39.12
TIES-Merging
41.85 - 0.0 20.93
DARE 41.47 - 33.00 37.24
DELLA 40.99 - 35.00 38.00
M3 (Ours) 45.20 - 35.20 40.2
Math
& Code
Average Merging
- 56.17 8.20 32.19
Task Arithmetic
- 64.67 8.60 36.64
TIES-Merging
- 64.67 22.60 43.64
DARE - 65.05 9.80 37.43
DELLA - 63.08 20.2 41.64
M3 (Ours) - 64.75 21.4 43.10
Appendix F Performance Drop Rate (PDR) for Adversarial Robustness
The adversarial robustness is evaluated using the Performance Drop Rate (PDR) (Zhu et al. 2023), which is defined as follows:
Appendix G Additional Experimental Results on Adversarial Robustness
@@
Appendix H Detailed Introduction to DARE
DARE (Drop And REscale) (Yu et al. 2024) is a model sparsification method designed to reduce the redundancy of delta parameters in fine-tuned models while preserving their task-specific capabilities. In SFT, model parameters are optimized to unlock abilities for specific tasks, with the difference between fine-tuned and pre-trained parameters referred to as delta parameters.
However, studies have shown that delta parameters are often highly redundant. DARE addresses this redundancy by randomly dropping a proportion $p$ of delta parameters (referred to as the drop rate) and rescaling the remaining ones by a factor of 1/(1p)1/(1-p). This simple yet effective approach enables DARE to eliminate up to 99% of delta parameters with minimal impact on model performance, particularly in large-scale models, and it can be applied using only CPUs.
Experiments on AlpacaEval, GSM8K, and MBPP reveal that the merged LM has the potential to outperform any individual source LM, presenting a significant new discovery. Notably, the 7B model obtained through DARE merging, SuperMario v2, ranks first among models of the same scale on the Open LLM Leaderboard (Beeching et al. 2023). These improvements were achieved without the need for retraining, positioning DARE as an efficient and resource-friendly solution for model merging.
Appendix I Integrating M3 into the TIES-Merging Model Merging Method